OpenStudy (anonymous):

In a 30°-60°-90° right triangle, the length of the shortest leg is s, what is the length of the hypotenuse?

OpenStudy (anonymous):

sin30=s/h h=s/sin30 h=s/.5 h=2s

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

do you understand what I did?

OpenStudy (anonymous):

yes i do